OpenStudy (anonymous):

A spherical balloon is being inflated at the rate of 8 cubic centimeters per second. At what rate is the radius changing when the volume = 10pi cc?

OpenStudy (anonymous):

so if im not mistaken this is very simple problem - normally \[\frac{ dv }{ dt }=4\pi r^2(\frac{ dr }{ dt })\]

OpenStudy (anonymous):

Yes, pretty simple, so now just plug the numbers in :)

OpenStudy (anonymous):

so if im trying to find out how fast the rate is changing then i need to move it around and solve for dr/dt \[\frac{ dr }{ dt }=\frac{ \frac{ dv }{ dt } }{ V }\] \[\frac{ dr }{ dt }=\frac{ 8 cc/\sec }{ 10 \pi cc } \]

OpenStudy (anonymous):

so is am i correct in saying that the radius is changing at 2.513 / sec ?

OpenStudy (anonymous):

if you dont mind all i did was just take 8/10pi and get 2.513?

OpenStudy (anonymous):

V' = 4π r² r' Now you need to find r² from V to plug in!

OpenStudy (anonymous):

-> r' = V' / ( 4π r² ) = ...

OpenStudy (anonymous):

arent i trying to find dr/dt? because i know that the V = 10pi ?

OpenStudy (anonymous):

but the ( 4π r² ) doesnt that = 10π in this equation?

OpenStudy (anonymous):

V' would = the 8 cc because that is the rate it is changing?

OpenStudy (anonymous):

You have V, but you need r² from V !

OpenStudy (anonymous):

super confused - what i think i know >> \[\frac{ dV }{ dt }=8 cc/\sec\] \[V = 10\pi cc\] what i need to know \[\frac{ dr }{ dt }\] \[\frac{ dr }{ dt } = \frac{ \frac{ dV }{ dt } }{ V }\]

OpenStudy (anonymous):

V' = 4π r² r' Do you agree up to this part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

-> r' = V' / ( 4π r² ) Do you understand this?

OpenStudy (anonymous):

yes i do :)

OpenStudy (anonymous):

so in this equation i should be trying to find what r² is, correct?

OpenStudy (anonymous):

Yes, sir, what we have is V' = -8, but we still miss r² to fill in

OpenStudy (anonymous):

since i dont know r' and i dont know r² i dont know how to solve the equation for r² :(

OpenStudy (anonymous):

Using the original V = (4/3 ) π r³ to find r : 3* 10π/ 4π = r³

OpenStudy (anonymous):

7.5π = r³ in this equation

OpenStudy (anonymous):

r² = 5.732

OpenStudy (anonymous):

r' = -8 / ( 4π (5.732 ) ?

OpenStudy (anonymous):

3* 10π/ 4π = r³ -> r³ = 7.5

OpenStudy (anonymous):

correct then r^2 would be 5.732

OpenStudy (anonymous):

r' = V' / ( 4π r² ) so if i plug it back into the r' equation i should get r' = -8 / ( 4π (5.732 ) correct?

OpenStudy (anonymous):

No, r² = 3.832

OpenStudy (anonymous):

sorry i feel dumb, pressing to many keys r' = -8 / ( 4π (3.832)

OpenStudy (anonymous):

So, you think you can handle from here :)

OpenStudy (anonymous):

Thank you, sorry for being dumb :)

OpenStudy (anonymous):

Take your time at start, then you'll be able to jump far in long run!